OpenStudy (anonymous):

2 + (x + y) = (2 + x) + y is an example of which algebraic property?

OpenStudy (magbak):

What are the options.

OpenStudy (anonymous):

Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (magbak):

It is the communicative property.

OpenStudy (anonymous):

This is the commutative property of addition. This basically says that the order in which you add numbers does not change the answer.

jimthompson5910 (jim_thompson5910):

it's not that

jimthompson5910 (jim_thompson5910):

commutative property is x+y = y+x or x*y = y*x

OpenStudy (anonymous):

And how is that different than what we have here @jim_thompson5910 ?

OpenStudy (magbak):

It is not the associative property because it is not x +(y+c)= y+(x+c)

OpenStudy (goldphenoix):

It is the associative propety. o-O

jimthompson5910 (jim_thompson5910):

it is the associative property because you're re-associating parenthesis

OpenStudy (uri):

I AM TELLING YOUU IT IS @magbak

OpenStudy (goldphenoix):

Yes. Your example fit the question's example, magbak. :)

OpenStudy (magbak):

No it does not.

OpenStudy (anonymous):

"In mathematics, a binary operation is commutative if changing the order of the operands does not change the result" Isn't that all we're doing here, changing the order in which we add things?

jimthompson5910 (jim_thompson5910):

the associative property of addition is x+(y+z) = (x+y)+z

OpenStudy (magbak):

OH YES IT DOES FIT THE EXAMPLE.

OpenStudy (goldphenoix):

Wait nevermind. Your example doesn't fit. But Associative property is the answer.

OpenStudy (magbak):

Yes it is the associative property of addition.

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (magbak):

You are welcome.

jimthompson5910 (jim_thompson5910):

vinnv226, because addition is commutative, this explains why associativity works but the two are different ideas

OpenStudy (anonymous):

Isn't the Associative Property just applied Commutative Property?

OpenStudy (magbak):

Yes @jim_thompson5910 is correct and should be given full credit.

OpenStudy (anonymous):

I'm still hung up on this and why the commutative property does not imply the exact same thing.

OpenStudy (goldphenoix):

Commutative Property = Switching the numbers and variable Associative Property = Switching the parenthesis only; the numbers and variable stays in the same place.

OpenStudy (anonymous):

But isn't the result the same? By switching the parenthesis, you're just switching the order in which you add things, which is the same as just switching the numbers around.

jimthompson5910 (jim_thompson5910):

if you're familiar with matrices and matrix multiplication you might recall that matrix multiplication is NOT commutative but it is associative ie A*B = B*A is false in general but A*(B*C) = (A*B)*C is true A,B,C are matrices

jimthompson5910 (jim_thompson5910):

so that's one example of how the two are different

OpenStudy (magbak):

Yes but computer wise and by checking by a automatically graded test it will be marked wrong unlike if it was a teacher graded assignment that a teacher can give half credit.

OpenStudy (anonymous):

But it seems to me that, if the Associative Property didn't exist, I could just as easily prove this statement by using the Commutative Property a few times.

jimthompson5910 (jim_thompson5910):

that's the thing: if an operation is commutative, then you can prove it's associative (not the other way around though)

jimthompson5910 (jim_thompson5910):

yes you can use the commutative property to say x+(y+z) = (x+y)+z is true, but the associative property is much quicker

OpenStudy (anonymous):

I see, that seems a little misleading then, to have both options as the answer on a test when both answers can be used to prove it. Seems equivalent to asking "which of these is a rectangle" and then having both a rectangle and a square as answer choices.

OpenStudy (anonymous):

Or rather, putting a picture of a square, asking "what shape is this" and having "square" and "rectangle" as answer choices

jimthompson5910 (jim_thompson5910):

if it's a square, then you would go with square since it's the most descriptive and closest to what it really is sure it's also a rectangle, but it's also a polygon too, but we don't go with "polygon" when we can say "square" since that's the most descriptive and to the point

OpenStudy (anonymous):

I see your point, but it strikes me as misleading. I guess thats why I don't write test questions.

jimthompson5910 (jim_thompson5910):

it is a bit misleading but it's just one of those questions where the best answer is chosen (more or less)

OpenStudy (anonymous):

Well, honestly it's uncommon I get to learn things about math while answering math questions on here, and even less common for me to learn something about a topic as basic as algebraic properties. Thanks for that Jim

jimthompson5910 (jim_thompson5910):

I'm glad I was of help

OpenStudy (magbak):

I was glade I was of no help and I just clarified so things.